# FedP2EFT: Federated Learning to Personalize Parameter Efficient Fine-Tuning for Multilingual LLMs

### v1

@@ A.3 Complementary Approaches and Base Models
Standard FL. Standard FL involves training a single global model. Given a pretrained LLM, we run FedAvg on the seen pool of clients, where $10\%$ of participating clients are sampled every round to train the model before sending the weights back for aggregation. In our XNLI and MasakhaNEWS setup, we do full fine-tuning of mBERT, setting each client’s learning rate to $55e-$ 5 and running FedAvg for 100 rounds. In our Fed-Aya setup, we adopt the training recipe from FedLLM-Bench (Ye et al., 2024) for MobileLLaMA-1.4B, where we do PEFT with LoRA applied to query and value attention weights ($r=16$, $\alpha_{lora}=32$, dropout$=0.05$) for 200 rounds. We use the cosine learning rate decay over 200 rounds with initial learning rate $2e^{-5}$ and minimum learning rate $1e^{-6}$.
Personalized FL. We train personalized base models using FedDPA-T (Yang et al., 2024b) and DEPT(SPEC) (Iacob et al., 2025) in our XNLI setup. FedDPA-T proposed having two separate LoRA adapters, one of which is shared (global) and the other is kept locally for each client (local). We adopted their training recipe for sequence classification, where the classifier is shared together with the global LoRA modules and the local LoRA modules stay local. LoRA modules are only applied to query and value attention weights (r=8, $\alpha_{lora}=8$, dropout=$0.05$). We set the learning rate to $5e^{-5}$.
@@ A.4 Baselines and FedP2EFT
LoRA PEFT (Hu et al., 2022). We search for the learning rate $[1e^{-5},1e^{-4},1e^{-3}]$ and the number of epochs $[1,2,3]$ and find that the learning rate $1e^{-4}$ with $2$ epochs had the best performance on the train set. We fixed $\alpha_{lora}=2r$. To ensure a similar inference budget across baselines, we set the number of epochs to $2$ for all our experiments.
AdaLoRA (Zhang et al., 2023). Similarly to LoRA, we search for the learning rate $[1e^{-5},1e^{-4},1e^{-3},1e^{-2}]$, the time interval between two budget allocations, $\Delta_{T}$, $[1,10,100]$ and the coefficient of the orthogonal regularization, $\gamma$, $[0.1,0.5]$. Within our search space, we find learning rate$=1e^{-3}$, $\Delta_{T}=1.0$, and $\gamma=0.1$ to be the best performing one. We run AdaLoRA once per resource budget $r$, setting the initial rank to be 1.5×1.5$\times$1.5 × of $r$, as recommended. We set the initial fine-tuning warmup steps and final fine-tuning steps to be $10\%$ and $30\%$ of the total steps respectively.
@@
BayesTune-LoRA (Section3.2). For fair comparison with FedP2EFT, we use the same hyperparameters as FedP2EFT. This baseline, hence, is an ablation study of how much performance collaboratively learning a PSG adds.
@@
FedL2P (Lee et al., 2023) As FedL2P requires a validation set for outer-loop bi-level optimization and federated early stopping, we split the train set of every client $80\%$ train and $20\%$ validation. Following FedL2P, we set the federated early stopping patience to 50 rounds, MLP hidden dimension is set to 100, the inner-loop learning rate to be the same as finetuning, $1e^{-4}$, and the hypergradient hyperparameters, $Q=3,\psi=0.1$ with hypergradient clamping of $[-1,1]$. We use Adam for both the inner-loop and outer-loop optimizers and search for the learning rate for the MLP (LRNet) $[1e^{-5},1e^{-4},1e^{-3}]$ and the learnable post-multiplier learning rate $\tilde{\eta}$ $[1e^{-4},1e^{-5},1e^{-6}]$, picking $1e^{-4}$ and $1e^{-6}$ to be the best respectively. Finally, we use $3$ outer-loop steps with an effective outer-loop batch size of $16$.
@@
FedP2EFT (Section 3.3) We set $\alpha_{r\_mul}=2$, our resulting $r_{init}$ is, hence, $32$ since our $r_{\text{max target}}=16$ for all experiments. Following our standard LoRA fine-tuning baseline, we adopt the same learning rate and $\alpha_{lora}$, $41e-$ 4 and $2r_{init}$ respectively. The learning rate of $\bm{\lambda}$, on the other hand, is searched $[1e^{-1},1e^{-2},1e^{-3},1e^{-4}]$, and we pick $1e^{-2}$ for all experiments. All $\lambda$ values are initialized to $1e^{-4}$. The MLP hidden dimension is set to 2×2$\times$2 × the input dimension, which is model dependent. We clamp the output of the MLP as well as $\lambda$ with a minimum value of $41e-$ 4 in the forward pass during training. Following FedL2P, we use a straight-through estimator (Bengio et al., 2013) after clamping to propagate gradients. We initialize the weights of MLP with Xavier initialization (Glorot & Bengio, 2010) using the normal distribution with a gain value of $61e-$ 6 and the bias with a constant $41e-$ 4. Lastly, we set $\alpha_{s}=1e^{+2}$ and $\alpha_{p}=1e^{-2}$ for all experiments.
Appendix B Additional Results

### v2

@@ A.3 Complementary Approaches and Base Models
Standard FL. Standard FL involves training a single global model. Given a pretrained LLM, we run FedAvg on the seen pool of clients, where $10\%$ of participating clients are sampled every round to train the model before sending the weights back for aggregation. In our XNLI and MasakhaNEWS setup, we do full fine-tuning of mBERT, setting each client’s learning rate to 5e55e-5 and running FedAvg for 100 rounds. In our Fed-Aya setup, we adopt the training recipe from FedLLM-Bench (Ye et al. 2024) for MobileLLaMA-1.4B, where we do PEFT with LoRA applied to query and value attention weights ($r=16$, $\alpha_{lora}=32$, dropout=0.05=0.05) for 200 rounds. We use the cosine learning rate decay over 200 rounds with initial learning rate $2e^{-5}$ and minimum learning rate $1e^{-6}$.
Personalized FL. We train personalized base models using FedDPA-T (Yang et al. 2024b) and DEPT(SPEC) (Iacob et al. 2025) in our XNLI setup. FedDPA-T proposed having two separate LoRA adapters, one of which is shared (global) and the other is kept locally for each client (local). We adopted their training recipe for sequence classification, where the classifier is shared together with the global LoRA modules and the local LoRA modules stay local. LoRA modules are only applied to query and value attention weights (r=8, $\alpha_{lora}=8$, dropout=$0.05$). We set the learning rate to $5e^{-5}$.
@@ A.4 Baselines and FedP2EFT
LoRA PEFT (Hu et al. 2022). We search for the learning rate $1e^{-5},1e^{-4},1e^{-3}]$ and the number of epochs $[1,2,3]$ and find that the learning rate $1e^{-4}$ with $2$ epochs had the best performance on the train set. We fixed $\alpha_{lora}=2r$. To ensure a similar inference budget across baselines, we set the number of epochs to $2$ for all our experiments.
[…]
@@
BT-LoRA (Section3.2). For fair comparison with FedP2EFT, we use the same hyperparameters as FedP2EFT. This baseline, hence, is an ablation study of how much performance collaboratively learning a PSG adds.
@@
FedL2P (Lee et al. 2023). As FedL2P requires a validation set for outer-loop bi-level optimization and federated early stopping, we split the train set of every client $80\%$ train and $20\%$ validation. Following FedL2P, we set the federated early stopping patience to 50 rounds, MLP hidden dimension is set to 100, the inner-loop learning rate to be the same as finetuning, $1e^{-4}$, and the hypergradient hyperparameters, $Q=3,\psi=0.1$ with hypergradient clamping of [1,1][-1,1]. We use Adam for both the inner-loop and outer-loop optimizers and search for the learning rate for the MLP (LRNet) $1e^{-5},1e^{-4},1e^{-3}]$ and the learnable post-multiplier learning rate $\tilde{\eta}$ $1e^{-4},1e^{-5},1e^{-6}]$, picking $1e^{-4}$ and $1e^{-6}$ to be the best respectively. Finally, we use $3$ outer-loop steps with an effective outer-loop batch size of $16$.
@@
FedP2EFT (Section 3.3). We set $\alpha_{r\_mul}=2$, our resulting $r_{init}$ is, hence, $32$ since our $r_{\text{max target}}=16$ for all experiments. Following our standard LoRA fine-tuning baseline, we adopt the same learning rate and $\alpha_{lora}$, 1e41e-4 and $2r_{init}$ respectively. The learning rate of $\bm{\lambda}$, on the other hand, is searched $1e^{-1},1e^{-2},1e^{-3},1e^{-4}]$, and we pick $1e^{-2}$ for all experiments. All $\lambda$ values are initialized to $1e^{-4}$. The MLP hidden dimension is set to $2\times$ the input dimension, which is model dependent. We clamp the output of the MLP as well as $\lambda$ with a minimum value of 1e41e-4 in the forward pass during training. Following FedL2P, we use a straight-through estimator (Bengio, Léonard, and Courville 2013) after clamping to propagate gradients. We initialize the weights of MLP with Xavier initialization (Glorot and Bengio 2010) using the normal distribution with a gain value of 1e61e-6 and the bias with a constant 1e41e-4. Lastly, we set $\alpha_{s}=1e^{+2}$ and $\alpha_{p}=1e^{-2}$ for all experiments.
A.5 Experiments Compute Resource
We used Flower (Beutel et al. 2020) [^8] https://github.com/adap/flower, which uses Ray (Moritz et al. 2018) [^9] https://github.com/ray-project/ray for distributed training and inference, for all our experiments. Each experiment in our XNLI and MasakhaNEWS setup was run with 4 to 6 NVIDIA A40s GPUs, with each client running on 1 GPU at a time. Each FedP2EFT federated training run took less than an hour in total while FedL2P training run took a few hours for each rank budget. As accuracy is cheap to compute, each inference run only took a few minutes to evaluate on all clients. Our Fed-Aya setup with Llama-3.2-3B, on the other hand, took under a day to federatedly train FedP2EFT and 1-3 days to train FedL2P for each rank budget on 8 NVIDIA A100s 80GB. Due to the higher cost of auto-regressive generative runs compared to classification, each inference run took roughly 1-3 hours.
Appendix B LoRA Rank Cost and Adaptability to Heterogeneous Devices
LoRA Rank Cost. In our experiments, we consider LoRA rank budgets of $r=2,4,8,16$ where the total rank budget is $r\cdot L$. Our empirical results on an NVIDIA A40 show a relatively small increase of 0.3GB in memory and 2.4 seconds in runtime between r=2 and r=16. This seemingly minor overhead can have major practical implications on devices with limited resources, such as laptops. Specifically, a 0.3GB memory footprint can be prohibitive on GPUs like QC Adreno 618 with typically only around 4GB of VRAM. We also predict a significant latency increase on mobile devices, as they lack the processing power and specialized hardware of larger systems and commonly are forced to offload parameters from RAM to the excessively slow-access flash memory, underscoring the need for efficient LoRA rank selection in real-world deployments. In scenarios where resources become more abundant, the server can host multiple models of different sizes and simply switch to a larger base model instead of using a higher rank. Nonetheless, we emphasize that our focus is to generate personalized fine-tuning hyperparameters for any base model chosen by the user.
Heterogeneous Devices Compatibility. During federated training, we assume that all clients that partake in training (seen clients) have sufficient resources to train the full model and the max LoRA rank. Clients with fewer resources are assumed to be in the unseen pool. Extending the work to train the PSG over heterogeneous resources is interesting as it allows clients with fewer resources to partake in the training stage. However, it is out of scope and we leave it for future work. During inference, we show that our approach mostly outperforms existing baselines with varying LoRA rank budgets across different model sizes.
Appendix C On BayesTune
BayesTune (Kim and Hospedales 2023) is a sparse fine-tuning approach that leverages a Bayesian framework to identify and update only the most important parameters of the model. It introduces a per-parameter scaling coefficient, where the magnitude of this coefficient governs the degree to which the corresponding parameter is updated during fine-tuning. These scaling coefficients, along with the model parameters, are explored and sampled from their posterior distribution using SGLD. To encourage sparsity, the Laplace and Gamma priors are imposed on the model parameters and their scaling coefficients respectively as discussed in Section 3.2.
Motivation for Adapting BayesTune. We adapt BayesTune for 2 reasons: 1) Most existing LoRA rank selection methods require separate runs for each rank resource budget while adapting BayesTune enables us to train only once for all targeted rank budgets. This is particularly useful as we only need to run the costly process of federated training once per base model. 2) The original proposed parameter-wise BayesTune is shown to be highly effective across a wide range of tasks, outperforming many existing pruning approaches and PEFT methods.
However, as mentioned in Section 3.2, directly using BayesTune in our approach will not only incur an intractable cost but also be difficult to optimize as the output dimension of our MLP will be too big and the likelihood of getting stuck in a suboptimal local minimum is high. This limitation can also be observed in existing FL hypernetwork approaches (Shamsian et al. 2021) where the experiments were constrained to learning the personalized parameters of a tiny LeNet-5. Similarly, the authors of FedL2P (Lee et al. 2023) found that extending these hypernetwork approaches to ResNet-18 is costly and often struggle to converge. To make federated hyperparameter optimization feasible, we scale down this output space by learning optimal LoRA rank coefficients instead. While doing so improves convergence stability, the output dimension scales linearly with the number of LoRA ranks, thereby also susceptible to the aforementioned limitation at higher ranks.
Besides limiting the output dimensionality through learning LoRA rank coefficients, we also opt to find a point estimate rather than a full characterization of the posterior distribution due to the limited number of data samples often found in cross-device FL. Doing so allows us to reduce the number of hyperparameters, namely SGLD’s noise schedule and level, to tune and simplify the optimization process, focusing valuable on-device computational resources on effectively learning the client-specific personalization strategy with the available limited data.
Choice of $\alpha_{g}$ and $\beta_{g}$. Following BayesTune, we aim to yield sparse models by putting high density on small $\lambda$ close to 0, penalizing $\lambda$ that deviates from 0. To achieve this, the Gamma prior imposed on $\lambda$ needs to meet several constraints: 1) it has mode close to 0 ($\alpha_{g}<1$ for mode=0) and, 2) it decays rapidly as $\lambda$ grows (mean of Gamma, $\frac{\alpha_{g}}{\beta_{g}}\approx 0$ and variance of Gamma, $\frac{\alpha_{g}}{\beta_{g}^{2}}\approx 0$). The choice $\alpha_{g}=0.01$ and $\beta_{g}=100$ is good enough to meet these constraints. In the original BayesTune work, the authors found there isn’t substantial difference in the final performance as long as $\alpha_{g}$, and $\beta_{g}$ satisfies the above constraints.
Appendix D FedP2EFT Algorithms
Complementing the federated training and inference pipeline (Figure 2 and 1, respectively), this section details Algorithm 1 (federated training) and Algorithm 2 (inference).
Algorithm 1 FedP2EFT: Federated Training
Input: A base model $\Phi^{i}_{BM}$ with $L$ layers for each client $i$.
1: initialize MLP $\phi$
2:for each federated round do
3: Random sample clients $\tilde{C}$
4:for client $i\in\tilde{C}$ do
5:  send $\phi$ to client
6:  Forward pass of local dataset to compute $E(h_{l}),SD(h_{l})$ for $1\leq l\leq L-1$
7:  Compute $\bm{\hat{\lambda}^{i,0}}$ using $\phi$ (Eq. 3.3)
8:  Initialize BayesTune-LoRA: Randomly initialize LoRA modules $\bm{A^{i}}\bm{B^{i}}$ and plug $\bm{\hat{\lambda}^{i,0}}$ in them
9:  $\bm{\hat{\lambda}^{i,s}}\leftarrow$ Finetune $\bm{A^{i}},\bm{\hat{\lambda}^{i,0}},\bm{B^{i}}$ for $s$ steps, minimizing the losses in Eq. 1
10:  $\phi^{i}\leftarrow$ Finetune $\phi$, minimizing L1 loss $\mathcal{L}_{1}(\bm{\hat{\lambda}^{i,0}},\bm{\hat{\lambda}^{i,s}})$
11:  send $\phi^{i}$ and num of data samples NN to server
12:end for
[…]
14:end for
Output: $\phi$
Algorithm 2 FedP2EFT: Inference
Input: A base model $\Phi^{i}_{BM}$ with $L$ layers for each client $i$. Learned MLP $\phi$.
1:for each client $i$ do
2: send $\phi$ to client
3: Forward pass of local dataset to compute $E(h_{l}),SD(h_{l})$ for $1\leq l\leq L-1$
4: Compute $\bm{\hat{\lambda}^{i}}$ using $\phi$ (Eq. 3.3)
[…]
6: Initialize BayesTune-LoRA for all layers $\lambda_{l}||_{1}>0$
7: Finetune BayesTune-LoRA layers $\bm{A^{i}},\bm{B^{i}}$ with frozen $\bm{\hat{\lambda}^{i}}$ and obtain personalized model $\Phi^{i}$
8:end for
Output: $\Phi^{i}$
Appendix E Additional Results